# Super-Resolution works for coastal simulations

### v1

@@ IV-C Ablation studies and analysis
$\bullet$ Losses for optimization. We propose to use three loss terms to optimize the proposed model. To show their effects, we visualize their effect by plotting the quantitative metrics in Figure 8. Given LR input at $t_{n}$ and $t_{n+1}$ and SR predictions at $t_{2n}$, $t_{2n+1}$, and $t_{2n+2}$, we take the coastal simulations at $t_{2n}$ and $t_{2n+2}$ (we refer to them as intra frames) to compute the spatial differences, and the coastal simulation at $t_{2n+1}$ (we refer to them as inter frames) to compute the temporal differences. We can see that using $L_{lp}$ and $L_{diff}$ losses can reduce both RMSE and MAE losses. Specifically, using $L_{diff}$ can significantly improve the temporal performance, approximately 24% to 42% in RMSE. It indicates that physics-informed losses can better supervise the temporal consistency for smooth solution changes.
@@
$\bullet$ Key modules for spatiotemporal matching. There are two key components in our proposed DNCSR : spatiotemporal attention and feature split and reconstruction (FSR). To illustrate their effects, we use the Bahamas dataset to conduct the ablation studies and report the RMSE, MAE, and SSIM in Table IV. As introduced in Figure 2, the feature extraction module can be any image feature encoders. We choose RDN [63], RCAN [73], and SwinIR [33] as candidates for the feature extraction. Then, we add spatiotemporal attention and/or feature split and reconstruction (FSR) to see if they can positively affect the super-resolution performance. We can see that 1) using SwinIR and RCAN is better than RDN as the feature encoder, as they can provide deeper feature representation, and rows 4 and 6 also prove that combining RCAN and spatiotemporal attention and FSR can provide the best performances. 2) Comparing rows 1&3 and 2&4, we can see that spatiotemporal attention can significantly improve the RMSE by about 0.021. Different feature encoders only have 0.08 improvements in RMSE and 0.002 in MAE. Given that SwinIR is a complex and time-consuming model, we choose RCAN, which balances super-resolution quality and computation efficiency well.
@@
TABLE IV: Ablation analysis on the key modules. We report the results on the test datasets of the Bahamas using different combinations of modules.
Module Evaluation
Feat encoder Spatiotemporal attn FSR RMSE MAE SSIM
RDN 0.054 0.023 0.971
RCAN 0.050 0.021 0.972
SwinIR 0.035 0.016 0.981
RDN 0.033 0.013 0.980
RCAN 0.031 0.013 0.981
SwinIR 0.025 0.011 0.982
RDN 0.030 0.012 0.982
RCAN 0.028 0.011 0.981
SwinIR 0.024 0.010 0.982
$\bullet$ Spatiotemporal attention. The key module of our proposed DNCSR is the spatiotemporal attention. To visually understand its ability to capture the pixel information for simulation, we utilize Local Attribution Maps (LAM) [79] to visualize the influences of neighborhood pixels on the region of interest (ROI) in Figure 9. The idea is to calculate the path-integrated gradient along a gradually changing path from the SR result to the LR input. In Figure 9, we show the contribution’s pixel attribute map and region in both $t_{n}$ and $t_{n+1}$ LR inputs. The red box region is calculated and enlarged for comparison. We can see that EDVR and ours can explore wider regions, which means that they can extract a wider range of correlated features for calculation. RCAN can also explore larger regions because of its channel attention computation, however, it fails to show the differences caused by the temporal changes. On the contrary, we can observe that ours shows attribution changes (the red color ranges become larger from $t_{n}$ to $t_{n+1}$ LR image) because ours can explore the temporal changes and reflect the attention score on the different LR images.
@@
Figure 9: Interpretation of SR method using attribute maps. The LAM maps represent the importance of each pixel in the input LR image w.r.t. the SR of the patch marked with a red box. We also show the area of contribution to highlight the ROI region for computing feature correlations.
Meanwhile, as introduced in Section III, the proposed spatiotemporal attention explores the horizontal(H)-, vertical(V)- and depth(D)-dimensional feature extraction hence we conduct the experiments to compare the combination of three sub-attention modules and report the results in Figure 10. On the top, we show the RMSE values of using attention operation at different dimensions. “No attention” means no attention is used in the network. “H+V+D+Pos” means our final model with the positional encoded spatiotemporal attention module. We can conclude that 1) combining horizontal, vertical, and depth attention achieves the best performance, and 2) from the feature maps, we can observe that the module can explicitly capture the dominant feature in corresponding directions.
@@
Figure 10: Visualization of spatiotemporal attention at different dimensions. On the top, we report the RMSE results of using different horizontal, vertical, and depth attention combinations. After applying horizontal, vertical, and depth attention, we show the feature map at the bottom.
$\bullet$ Feature split and reconstruction. For reconstruction, we are inspired by the recent development of image processing in the frequency domain [80]. We apply the Fast Fourier Transform (FFT) to the LR image and the feature maps to see if they learn different aspects of feature representation. As shown in Figure 11, we show the feature map after cosine and sine operations. The magnitude spectrum maps show that they learn complementary information to predict the SR result. For instance, the cosine feature map learns the global low-frequency information, and the spectrum energy concentrates in the center. The sine feature map learns more about edge information, and the spectrum energy focuses on the vertical line.
@@
$\bullet$ Motion consistency. We calculate the RMSE score of the Bahamas dataset to measure the motion consistency. We report intra-frame and inter-frame super-resolution results to see if the RMSE score changes due to the spatial and temporal upsampling. We do not use optical flow or other motion estimation models because they are built on key feature point tracking. Our experiments show that they fail to estimate the dynamics of the coastal simulation. Hence, we measure separately for intra-frame and inter-frame super-resolution to calculate the spatial and temporal RMSE losses and plot them in Figure 12. We can see that our results on spatial and temporal losses are almost identical. At the same time, EDVR and RDN have big disagreements on the spatial and temporal super-resolution. It indicates that ours can consistently super-resolve intra- and inter-frames with good quality without compromising the motion changes.
Figure 12: Plot of frame-by-frame RMSE scores by using different methods. The temporal and spatial RMSE values for the Bahamas dataset validating the model performance on motion consistency.
V Conclusion
@@
In this paper, we propose DNCSR , a spatiotemporal super-resolution neural network for efficient coastal ocean simulation. The proposed spatiotemporal attention fully utilizes the neighborhood pixels across space and time via pixel shuffle to estimate the feature correlations. It can explicitly model the U-, V- and $\xi$-view changes over time as spatial feature correlations. The grid location is also embedded as the positional code for axis-aware attention estimation. Finally, we implement the novel feature split and reconstruction to project the features into the frequency domain for final enhancement. The initial spatiotemporal bilinear operator is used to encourage the network to learn the residues between LR and HR data. The proposed physics-informed losses also help to optimize the neural network with better quantitative results. In order to train the neural network, we propose large-scale coastal simulation datasets on the Bahamas and Galveston. We use it for model training and evaluation. From extensive experiments and comparisons with other state-of-the-art methods, we can conclude that ours can achieve superior super-resolution quality with fast computation.
@@
Since the proposed methodology guarantees the mass conservation for arbitrarily long simulation times, is is particularly attractive for ocean, climate, and weather models for which high resolution simulations are desirable but computationally very expensive. In addition, it paves a novel direction for neural super-resolution models for multigrid simulation, and we are interested in further exploration of ultra-resolution reconstruction and continuous super-resolution.
@@ Acknowledgments
A. Rupp has been supported by the Academy of Finland’s grant number 350101 Mathematical models and numerical methods for water management in soils, grant number 354489 Uncertainty quantification for PDEs on hypergraphs, grant number 359633 Localized orthogonal decomposition for high-order, hybrid finite elements, Business Finland’s project number 539/31/2023 3D-Cure: 3D printing for personalized medicine and customized drug delivery, and the Finnish Flagship of advanced mathematics for sensing, imaging and modeling, decision number 358944.

### v2

@@ 5.3 Ablation studies and analysis
We propose to use three loss terms to optimize the proposed model. To show their effects, we visualize their effect by plotting the quantitative metrics in Figure 8. Given LR input at $t_{n}$ and $t_{n+1}$ and the predictions at $t_{2n}$, $t_{2n+1}$, and $t_{2n+2}$, we take the coastal simulations at $t_{2n}$ and $t_{2n+2}$ (we refer to them as intra frames) to compute the spatial differences, and the coastal simulation at $t_{2n+1}$ (we refer to them as inter frames) to compute the temporal differences. We can see that using $L_{lp}$ and $L_{diff}$ losses can reduce both RMSE and MAE losses. Specifically, using $L_{diff}$ can significantly improve the temporal performance, approximately 24% to 42% in RMSE. It indicates that physics-informed losses can better supervise the temporal consistency for smooth solution changes.
@@
There are two key components in our proposed DNNCS : spatiotemporal attention (ST Attn) and feature split and reconstruction (FSR). To illustrate their effects, we use the Bahamas dataset to conduct the ablation studies and report the RMSE, MAE, and SSIM in Table 4. As introduced in Figure 2, the feature extraction module can be any image feature encoders. We choose RDN [85], RCAN [84], and SwinIR [50] as candidates for the feature extraction. Then, we add spatiotemporal attention and/or feature split and reconstruction (FSR) to see if they can positively affect the downscaling performance. We can see that 1) using SwinIR and RCAN is better than RDN as the feature encoder, as they can provide deeper feature representation, and rows 4 and 6 also prove that combining RCAN and spatiotemporal attention and FSR can provide the best performances. 2) Comparing columns 1&4 and 2&5, we can see that spatiotemporal attention can significantly improve the RMSE by about 0.021. Different feature encoders only have 0.08 improvements in RMSE and 0.002 in MAE. Given that SwinIR is a complex and time-consuming model, we choose RCAN, which balances downscaling quality and computation efficiency well.
@@
Table 4: Ablation analysis on the key modules. We report the results on the test datasets of the Bahamas using different combinations of modules. ST Attn refers to Spatiotemporal attention operation.
Module Feat encoder RDN RCAN SwinIR RDN RCAN SwinIR RDN RCAN SwinIR
ST Attn
FSR
Evaluation RMSE 0.054 0.050 0.035 0.033 0.031 0.025 0.030 0.028 0.024
MAE 0.023 0.021 0.016 0.013 0.013 0.011 0.012 0.011 0.010
SSIM 0.971 0.972 0.981 0.980 0.981 0.982 0.982 0.981 0.982
@@
Figure 9: Interpretation of downscaling method using attribute maps. The LAM maps represent the importance of each pixel in the input LR image w.r.t. the downscaled results of the patch marked with a red box. We also show the area of contribution to highlight the ROI region for computing feature correlations.
@@
Figure 10: Visualization of spatiotemporal attention at different dimensions. On top, we report the RMSE results of using different horizontal, vertical, and depth attention combinations. After applying horizontal, vertical, and depth attention, we show the feature map at the bottom.
Spatiotemporal attention
The key module of our proposed DNNCS is the spatiotemporal attention. To visually understand its ability to capture the pixel information for simulation, we use Local Attribution Maps (LAM) [35] to visualize the influences of neighborhood pixels on the region of interest (ROI) in Figure 9. The idea is to calculate the path-integrated gradient along a gradually changing path from the downscalied result to the LR input. In Figure 9, we show the contribution’s pixel attribute map and region in both $t_{n}$ and $t_{n+1}$ LR inputs. The red box region is calculated and enlarged for comparison. We can see that EDVR and ours can explore wider regions, which means that they can extract a wider range of correlated features for calculation. RCAN can also explore larger regions because of its channel attention computation, however, it fails to show the differences caused by the temporal changes. On the contrary, we can observe that ours shows attribution changes (the red color ranges become larger from $t_{n}$ to $t_{n+1}$ LR image) because ours can explore the temporal changes and reflect the attention score on the different LR images.
@@
Figure 12: Plot of frame-by-frame RMSE scores by using different methods. It compares RMSE across 165 frames for different methods, evaluating spatial and temporal consistency. Temporal RMSE curves for RDN, EDVR and PhySR show prominent oscillations and higher overall error, indicating instability and poor motion coherence across frames. In contrast, spatial RMSE remains relatively flat, suggesting more consistent intra-frame quality. Our method (in pink for temporal and red for spatial) shows rapid convergence within the first 20 frames and maintains the lowest and most stable RMSE afterward, especially in the temporal domain, demonstrating superior consistency over time. There are 8 lines. It is just that some lines are very close to each other. I added some description to the caption and main text.
Meanwhile, as introduced in Section III, the proposed spatiotemporal attention explores the horizontal(H)-, vertical(V)- and depth(D)-dimensional feature extraction hence, we conduct the experiments to compare the combination of three sub-attention modules and report the results in Figure 10. On top, we show the RMSE values of using attention operation at different dimensions. “No attention” means no attention is used in the network. “H+V+D+Pos” means our final model with the positional encoded spatiotemporal attention module. We can conclude that 1) combining horizontal, vertical, and depth attention achieves the best performance, and 2) from the feature maps, we can observe that the module can explicitly capture the dominant feature in corresponding directions.
Feature split and reconstruction
For reconstruction, we are inspired by the recent development of image processing in the frequency domain [48]. We apply the Fast Fourier Transform (FFT) to the LR image and the feature maps to see if they learn different aspects of feature representation. As shown in Figure 11, we show the feature map after cosine and sine operations. The magnitude spectrum maps show that they learn complementary information to predict the downscaled result. For instance, the cosine feature map learns the global low-frequency information, and the spectrum energy concentrates in the center. The sine feature map learns more about edge information, and the spectrum energy focuses on the vertical line.
Motion consistency
We calculate the RMSE score of the Bahamas dataset to measure the motion consistency. We report intra-frame and inter-frame downscaling results to see if the RMSE score changes due to the spatial and temporal upsampling. We do not use optical flow or other motion estimation models because they are built on key feature point tracking. The figure illustrates RMSE trends over 165 frames for different super-resolution models. Our method (magenta for temporal, red for spatial) exhibits nearly identical curves in both spatial and temporal cases, indicating high motion consistency and robust reconstruction across frames. In contrast, RDN (blue and green) shows a large discrepancy between spatial and temporal RMSE, with noticeable oscillations in the temporal curve, revealing poor temporal coherence. EDVR and PhySR demonstrate smaller spatial-temporal mismatches (the gray and yellow curves of PhySR and light and dark brown curves of EDVR), but both produce higher RMSE overall compared to ours. These results highlight that our method achieves the best balance between spatial fidelity and temporal stability, making it particularly suitable for dynamic scenes like coastal simulations where motion preservation is critical.
5.4 Limitations and further discussion
One potential shortcoming of our proposed model is that it should be fine-tuned on other unknown regions for better performance. This is due to the domain-specific physical conditions at different places. For instance, the shoreline can affect fine-grained water movements. In our design, we define the landmass as zero, which serves as a hard condition for the downscaling neural network. One promising approach is to convert the coastal simulation into graphs and use graph neural networks to learn downscaling as graph interpolation. The advantage is that the water movement can pass along the vertices and edges and find nonlocal correlations for downscaling.
6 Conclusion
@@
In this paper, we propose DNNCS , a spatiotemporal downscaling neural network for efficient coastal ocean simulation. The proposed spatiotemporal attention fully utilizes the neighborhood pixels across space and time via pixel shuffle to estimate the feature correlations. It can explicitly model the U-, V- and $\xi$-view changes over time as spatial feature correlations. The grid location is also embedded as the positional code for axis-aware attention estimation. Finally, we implement the novel feature split and reconstruction to project the features into the frequency domain for final enhancement. The initial spatiotemporal bilinear operator is used to encourage the network to learn the residues between LR and HR data. The proposed physics-informed losses also help to optimize the neural network with better quantitative results. In order to train the neural network, we propose a novel coastal simulation dataset on the Bahamas and Galveston. We use it for model training and evaluation. From extensive experiments and comparisons with other state-of-the-art methods, we can conclude that ours can achieve superior downscaling quality with fast computation.
@@
Since the proposed methodology guarantees the mass conservation for arbitrarily long simulation times, it is particularly attractive for ocean, climate, and weather models for which high resolution simulations are desirable but computationally very expensive. In addition, it paves a novel direction for neural downscaling models for multigrid simulation, and we are interested in further exploration of ultra-resolution reconstruction and continuous downscaling.
@@ Acknowledgments
A. Rupp has been supported by the Academy of Finland’s grant number 350101 Mathematical models and numerical methods for water management in soils, grant number 354489 Uncertainty quantification for PDEs on hypergraphs, grant number 359633 Localized orthogonal decomposition for high-order, hybrid finite elements, Business Finland’s project number 539/31/2023 3D-Cure: 3D printing for personalized medicine and customized drug delivery, and the Finnish Flagship of advanced mathematics for sensing, imaging and modeling, decision number 358944. Markus Büttner was partially supported by the German Research Foundation (DFG) under grant AI 117/7-1 within the project 502500606 ’Performance-optimized co-design of ocean modeling software on FPGAs’.